OpenStudy (anonymous):

(9-4i)-(9-4i)=91-72i???

OpenStudy (anonymous):

81-72i is the answer??

OpenStudy (hunus):

The answer would be zero

OpenStudy (anonymous):

zero.....? that is definitly not what i got

OpenStudy (anonymous):

You're basically subtracting one thing with the same thing due to the parenthesis so it's 0

OpenStudy (hunus):

You're subtracting (9-4i) from itself

OpenStudy (anonymous):

the original question was (9-4i)^2

OpenStudy (hunus):

Oh

OpenStudy (anonymous):

ohh in that case you would foil

OpenStudy (anonymous):

i dont think the answer is zero

OpenStudy (anonymous):

so i got 81-72i just dont think my answer is right but i know it's not 0

OpenStudy (hunus):

\[(9-4i)^2=(9-4i)(9-4i)=9*9-4i*9-4i*9+4i*4i\]

OpenStudy (anonymous):

yes hunus that was my steps as well

OpenStudy (hunus):

The question your posted says (9-4i)-(9-4i)

OpenStudy (anonymous):

sorry i was trying to show the first step i did

OpenStudy (hunus):

It's fine :) first \[(9−4i)^2=(9−4i)(9−4i)=9∗9−4i∗9−4i∗9+4i∗4i\] then \[9∗9−4i∗9−4i∗9+4i∗4i = 81 -36i - 36i +16i^2 = 81 - 72i -16=65 - 72i\]

OpenStudy (anonymous):

i definitely found my mistake thank you so much.

OpenStudy (anonymous):

i gave you a medal

OpenStudy (hunus):

Yup :)

OpenStudy (hunus):

thanks :D